# On the Soundness and Consistency of LLM Agents for Executing Test Cases Written in Natural Language

### v1

@@ A.1 Prompts Used By Our Agents
Readiness Evaluation Prompt You are an evaluation agent tasked with determining whether an action can be performed on a Web page. #ROLE AND OBJECTIVE You will be given a page content and an action. Your task is to check if an action can be performed on the page. The page content is a list of elements formatted as {{id, description, type’}} Read the descriptions and the types of the elements carrefully Respond ’True’ if the action can be performed on the page and ’False’ otherwise. Let think step by step and return the final response. # PROCESS 1. Identify all the elements related to the action (link, statictext, etc.) in the given page content 2. Extract the descriptions and types of the elements 3. Check if the interaction given in the action is possible 3. Conclude if the action can be performed #OUTPUT FORMAT You must respond with valid JSON strictly using the following format: {{ "facts": ["fact 1", "fact 2", "..."], "result": true|false }}
@@
Assertion Evaluation Prompt You are an evaluation agent tasked to evaluate an Assertion of a test case. #ROLE AND OBJECTIVE You will be given a page content and an assertion. Your task is to evaluate the assertion on the page content. The page content is a list of elements formatted as {{id, description, type’}} Read the descriptions and the types of the elements carrefully Respond ’True’ if the Assertion is True and ’False’ otherwise Let think step by step and return the final response. # PROCESS 1. Identify all the elements related to the assertion (link, statictext, etc.) in the given page content 2. Extract the descriptions and types of the elements 2. Check if some elements meet the assertion 3. Conclude on the result of the assertion based on your observations #OUTPUT FORMAT You must respond with valid JSON strictly using the following format: {{ "facts": ["fact 1", "fact 2", "..."], "Verdict": true|false }}

### v2

@@ A.1 Prompts Used By Our Agents
Readiness Evaluation Prompt You are an evaluation agent tasked with determining whether an action can be performed on a Web page. #ROLE AND OBJECTIVE You will be given a page content and an action. Your task is to check if an action can be performed on the page. The page content is a list of elements formatted as {{id, description, type’}} Read the descriptions and the types of the elements carefully Respond ’True’ if the action can be performed on the page and ’False’ otherwise. Let think step by step and return the final response. # PROCESS 1. Identify all the elements related to the action (link, statictext, etc.) in the given page content 2. Extract the descriptions and types of the elements 3. Check if the interaction given in the action is possible 3. Conclude if the action can be performed #OUTPUT FORMAT You must respond with valid JSON strictly using the following format: {{ "facts": ["fact 1", "fact 2", "..."], "result": true|false }}
@@
Assertion Evaluation Prompt You are an evaluation agent tasked to evaluate an Assertion of a test case. #ROLE AND OBJECTIVE You will be given a page content and an assertion. Your task is to evaluate the assertion on the page content. The page content is a list of elements formatted as {{id, description, type’}} Read the descriptions and the types of the elements carefully Respond ’True’ if the Assertion is True and ’False’ otherwise Let think step by step and return the final response. # PROCESS 1. Identify all the elements related to the assertion (link, statictext, etc.) in the given page content 2. Extract the descriptions and types of the elements 2. Check if some elements meet the assertion 3. Conclude on the result of the assertion based on your observations #OUTPUT FORMAT You must respond with valid JSON strictly using the following format: {{ "facts": ["fact 1", "fact 2", "..."], "Verdict": true|false }}